OpenStudy (anonymous):

find the value of x. |x-4|=5

OpenStudy (anonymous):

-1, 9

OpenStudy (saifoo.khan):

| x - 4 | = 5 or | x - 4 | = -5

OpenStudy (saifoo.khan):

solve both!

OpenStudy (anonymous):

9

OpenStudy (mimi_x3):

x - 4 = 5 x = 9 x - 4 = -5 x = -1

OpenStudy (saifoo.khan):

x - 4 = 5 or x - 4 = -5

OpenStudy (anonymous):

got it thanks